For the reaction H2(g) + Cl2(g) 2 HCl(g), the equilibrium constant K at 800oC is 4.35 x 10^4.

Hydrogen at a partial pressure of 0.700 bar and chlorine at a partial pressure of 0.500 bar are placed in a vessel at 800oC and allowed to equilibrate. Find the final partial pressures of all three gases in this reaction in bar
Answer #1

-

------H2(g)   +    Cl2(g) 2 HCl(g)

I ----- 0.7 ----       0.5 --------------   0

C----- -x ----------- -x ----------------- +2x

E --- 0.7-x ------ 0.5-x ------------ 2x

     Kp    = P^2HCl/PH2*PCl2

    4.35*10^4   = (2x)^2/(0.7-x)(0.5-x)

    4.35*10^4(0.7-x)(0.5-x) = 4x^2

      x = 0.499

PH2   = 0.7-x =0.7- 0.499 = 0.201bar

PCl2   =0.5-x =0.5- 0.499 = 0.001bar

PHCl = 2x    = 2*0.499 bar = 0.998bar